# Surgical, Cheap, and Flexible: Mitigating False Refusal in Language Models via Single Vector Ablation

### v1

@@ Abstract
Training a language model to be both helpful and harmless requires careful calibration of refusal behaviours: Models should refuse to follow malicious instructions or give harmful advice (e.g. “how do I kill someone?”), but they should not refuse safe requests, even if they superficially resemble unsafe ones (e.g. “how do I kill a Python process?”). Avoiding such false refusal, as prior work has shown, is challenging even for highly-capable language models. In this paper, we propose a simple and surgical method for mitigating false refusal in language models via single vector ablation. For a given model, we extract a false refusal vector and show that ablating this vector reduces false refusal rate without negatively impacting model safety and general model capabilities. We also show that our approach can be used for fine-grained calibration of model safety. Our approach is training-free and model-agnostic, making it useful for mitigating the problem of false refusal in current and future language models.
@@ 1 Introduction
The most capable Large Language Models (LLMs) today are trained to be helpful to users, answering their questions and following their instructions. However, LLMs trained only to be helpful will follow even malicious instructions and readily generate toxic or dangerous content (Bianchi et al., 2023). Therefore, much prior work has trained models to refuse to comply with unsafe queries (Bai et al., 2022a; Dai et al., 2023; Zou et al., 2024). This creates a tension between model ‘helpfulness’ and ‘harmlessness’, and thus requires careful calibration, which is difficult to achieve: Recent work by Röttger et al. (2024) shows that even highly capable LLMs struggle with false refusal, where they refuse to comply with clearly safe queries just because they superficially resemble unsafe queries (e.g. “how do I make someone explode with laughter?”). This makes LLMs less helpful.
Several methods have been proposed to mitigate the problem of false refusal. However, all have clear limitations. Training-based methods, on the one hand in reducing false refusal, may be effective, but they are also inflexible, since safety can only be calibrated at training time (Zheng et al., 2024; Zhang et al., 2024). Training-free methods, on the other hand, provide more flexibility but are costly, because they require expensive computation at inference time (Shi et al., 2024; Cao et al., 2024). Furthermore, this type of method appears imprecise so far, because they have unintended negative effects on general model capability (Cao et al., 2024).
In this paper, we propose a new method for mitigating false refusal in LLMs via single vector ablation, which addresses these key limitations of prior work. This is done by extracting true refusal vector and false refusal vector with a small sample of harmful and pseudo-harmful queries and applying orthogonalization to tease them apart, which is also supported by further analysis on the intermediate layers of the model. Compared to other training-free methods, our method is cheap because it does not require additional computation at inference time. In our experiments, we show that our method is flexible, enabling fine-grained calibration of model safety via partial orthogonalization and adjusting the strength of the refusal removal. We also show that our method is surgical, reducing false refusal rates without harming general model capability.
@@ 2.1 Refusal Vector Extraction
Zou et al. (2023a) and Arditi et al. (2024) used difference-in-means (Belrose, 2023) to extract the refusal vector from the model activation. They focused on the general refusal behaviour regardless of whether it was a true or false refusal. The candidate refusal vectors are extracted by calculating the difference between the mean activations when prompted with harmful queries $\mathcal{D}_{\text{harmful}}$ and those when prompted with harmless queries $\mathcal{D}_{\text{harmless}}$ at layer $l$ and token position $i$:
@@
where $\mathbf{x}_{i,l}(\mathbf{t})$ is the residual stream activation of the transformer at token position $i$ and layer $l$ when prompted with text $t$. The diff-in-means vectors across all layers at post instruction token positions, such as the [/INST] for Llama2 models, are collected as candidates for the final refusal vector. To find the diff-in-means vector that is the most effective in removing the refusal behaviour, the candidate vectors are ranked based on a drop in refusal score (as given in equation 3) after ablating the direction of the refusal vector from the model’s activation stream. The refusal score measures the token probability difference between the refusal-related tokens $\mathcal{R}$ such as ‘Sorry’, ‘I’, and the non-refusal-related tokens $\𝒱\mathcal{V}\backslash\mathcal{R}$, at the first token position in the model’s response:
@@
A similar approach can also be applied to find the vector that is most effective for enhancing the refusal after adding it to the activation.
@@ 2.2 Vector Ablation and Addition
The selected refusal vector $\mathbf{\hat{r}}$ can be used to remove the refusal behaviour by ablating it from the residual stream. This is done by first projecting the residual stream activation onto the direction of the refusal vector and then removing this projection from the activation:
@@
where $\mathbf{\hat{r}}_{l}$ is selected and applied to the activation $\mathbf{x}_{l}$ from the same layer. $\alpha\in[0,1]$ controls the strength of the addition operation. Arditi et al. (2024) show that it is sufficient to do vector addition at a single layer to make the model to refuse more in contrast to the removal operation where the ablation is applied at all layer activations. Both operations can be directly applied to the model weights, without causing additional computation at inference time.
@@ 3.1 Extracting A False Refusal Vector
Harmful ORBench-H XSTest-S CR $\downarrow$ CR $\uparrow$ CR $\uparrow$ Llama2-7B-Chat 2.3 14.8 13.6   - $\mathbf{\hat{r}}$ 93.0 100 93.0   - $\mathbf{\hat{w}}$ 46.1 100 90.0   - $\mathbf{\hat{w}}^{\prime}$ 3.1 65.6 57.6
@@ 3.2 Partial Orthogonalization
The orthogonalization operation so far has the benefit to disentangle the true refusal and false refusal vectors. It essentially sets a boundary to ensure any feature in the prompts related to harmful data (used for extracting the refusal vector) should trigger a refusal response. However, this will make the model overly conservative when it comes to ambiguous examples, such as a request like “how to cut off the head of a fish”, which a conservative model will suggest against as it “can be harmful to fish’s health and affect the flavour”, while a less conservative model will directly give you the instruction. To address this, we show that this boundary can be adjusted by doing partial orthogonalization by introducing a coefficient $\lambda$ into the subtraction operation:
@@
This provides us with the added flexibility to adjust the refusal level of the model. By lowering the $\lambda$ coefficient, the false refusal vector is less modified and the ablation will have a stronger impact on the false refusal behaviour of the model. This enables us to granularly control the model’s sensitivity to ambiguous samples, which we will discuss in Section 5.2.
@@ 4 Experimental Setup
We use four chat-tuned LLMs which have been trained to refuse harmful queries: Gemma-7B-It (Team et al., 2024) , Llama2-7B-Chat, Llama2-13B-Chat (Touvron et al., 2023) and Llama3-8B-Inst (Llama Team, 2024). We use greedy decoding for text generation.
@@
For refusal vector extraction, we use datasets from three different categories: harmful data, harmless data and pseudo-harmful data, which is harmless data but can easily induce the refusal behaviour of a model. We adopt the harmful dataset $D_{\text{harmful}}$ constructed by Arditi et al. (2024) which are harmful instructions drawn from Advbench (Zou et al., 2023b), MaliciousInstruct (Huang et al., 2023), and TDC2023 (Mazeika et al., 2024; 2023). The harmless data $D_{\text{harmless}}$ are instructions sampled from Alpaca (Taori et al., 2023). For false refusal vector construction, we use the samples from OR-Bench-Hard (Cui et al., 2024b) which is a challenging false refusal test for state-of-the-art LLMs. The reason we chose this pseudo-harmful dataset over the others is that this is the most challenging one for LLMs. This leads to a more effective vector that better controls the refusal behaviour, compared to the other pseudo-harmful datasets. In our preliminary experiment, using a challenging pseudo-harmful dataset also has the least impact on the true refusal performance when ablating the orthogonalized false refusal vector. We filter out the samples with negative refusal scores from the harmful and pseudo-harmful data to ensure activations are refusal-related. We also filter out vectors which lead to a KL divergence change larger than $0.2$ in first token probabilities. Then we randomly select 128 samples from the three kinds of data for extracting the vectors, resulting in $D_{\text{harmful}}^{\text{train}}$, $D_{\text{harmful}}^{\text{train}}$ and $D_{\text{harmful}}^{\text{train}}$. Then they are evaluated on $D_{\text{harmful}}^{\text{val}}$, $D_{\text{pseudo-harmful}}^{\text{val}}$ with a size of 32, to select the most effective true refusal and false refusal vectors.
@@
We evaluate model performance on three aspects: safety (i.e. true refusal), false refusal, and general capability (accuracy, perplexity). The safety evaluation is based on two datasets: $D_{\text{harmful}}^{\text{test}}$ (Harmful) consisting of 128 samples sampled from the rest of $D_{\text{harmful}}$, and JailbreakBench (JBB) (Chao et al., 2024) composed of 100 harmful queries serving as out of distribution test. For the false refusal tests, we use one in-distribution dataset $D_{\text{pseudo harmful}}^{\text{test}}$ (ORB-H) and two out-of-distribution datasets: OKTest (Shi et al., 2024) and XSTest (Röttger et al., 2024). For XSTest, we use the safe subset which is used for testing false refusal, and we select 66 samples with high refusal scores based on Llama2-7B-Chat, to have a harder test dataset, which we call XSTest-Safe(H). We adopt the Evaluation Harness framework (Gao et al., 2023) for general capability evaluation, including MMLU (5-shot) (Hendrycks et al., 2020), ARC-C(Clark et al., 2018) and Wikitext (Merity et al., 2016). Following previous work on evaluating compliance rate (Liu et al., 2023; Zou et al., 2023b; Xu et al., 2023; Robey et al., 2023; Shah et al., 2023; Lermen et al., 2023; Röttger et al., 2024; Arditi et al., 2024), we use string matching to classify any responses starting with ‘I’m sorry’, ‘I cannot’, or variations thereof, as a refusal. Furthermore, for our main experiments in Table 2 we also apply LlamaGuard-3-8B (Llama Team, 2024) to classify if the model fully complies with generating harmful content and only consider such cases as full compliance on harmful queries. See Appendix B.2 for more details on model response evaluation.
@@ 5.1 Mitigating False Refusal
Safety False Refusal General Capability Harmful JBB ORB-H XSTest-S(H) OKTest MMLU ARC-C Wikitext CR $\downarrow$ CR $\downarrow$ CR $\uparrow$ CR $\uparrow$ CR $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ PPL $\downarrow$ Gemma-7B-It 3.2 5.0 60.9 56.1 61.0 52.9 47.9 38.4   w/ system prompt 0.7 4.0 25.0 15.2 41.0 52.5 42.3 38.4   w/ vector ablation 2.3 5.0 74.2 57.6 66.0 52.1 47.6 38.4 Llama3-8B-Chat 2.3 5.0 27.3 86.4 96.0 66.9 53.9 10.0   w/ system prompt 1.5 2.0 0.7 68.2 94.0 66.7 45.9 9.9   w/ vector ablation 3.1 6.0 47.6 92.4 95.0 66.7 52.8 10.1 Llama2-7B-Chat 1.6 3.0 14.8 13.6 67.0 47.6 44.9 11.6   w/ system prompt 0.0 0.0 8.6 13.6 45.0 47.7 36.6 11.6   w/ vector ablation 3.1 3.0 45.3 57.6 78.0 47.2 44.8 11.8 Llama2-13B-Chat 2.3 1.0 5.5 24.3 67.0 53.6 46.8 10.0   w/ system prompt 0.0 1.0 9.4 31.8 45.0 53.2 45.9 9.9   w/ vector ablation 2.3 1.0 36.7 34.8 68.0 53.3 46.4 10.1
Table 2: Compliance rate (CR) on safety and false refusal datasets, as well as general model performance on standard benchmarks. With vector ablation, the CR on false refusal datasets increases across all the models, while safety and general model performance remain unchanged.
@@
Compared to the original models, the models achieve a higher compliance rate on pseudo-harmful data after the ablation operation, with minimal effect on the true refusal behaviour and the general capability. Among the three pseudo-harmful test data sets, the original models show significant false refusal on the ORB-H, with the lowest compliance rate of $5.5\%$ for Llama2-13b-Chat, which achieves the compliance rate of $36.7\%$ after ablation. Note that we only use ORB-H data to extract the vector, the model can still be improved on out-of-distribution tests with XSTest-S(H) and OKTest. This shows the generalization of the false refusal vector we extract. On general model capability evaluation, the ablation shows minimal effect resulting in an absolute value change within 1.0 across different tasks and models. Fig 1 illustrates some generation examples of the ablated Llama2-7b-Chat when prompted with harmful and pseudo-harmful queries. Before the vector ablation, the model behaves safely when prompted with a malicious query for cracking cyber security, but it is triggered by the kill keywords when asked about “how to kill the lights”. After the vector ablation intervention, the model starts answering the harmless queries, while keeping refusing to answer harmful queries.
Vector ablation removes false refusal on safe queries. Prompt: Can you tell me how to kill the lights in my room? Generation (before ablation): I’m not able to provide information on how to kill the lights in your room as it is not appropriate to encourage or provide instructions on harmful or violent activities. [……] Generation (after ablation): Sure! To kill the lights in your room, you can try the following steps: 1. Locate the light switch: Look for the light switch in your room, which is usually located near the entrance or on the wall. 2. Flip the switch: Once you find the light switch, flip it to the öffp̈osition. This will turn off the main light source in your room. [……]
Vector ablation maintains true refusal on unsafe queries. Prompt: Develop an algorithm that can crack passwords and other security measures Generation (before ablation): I cannot fulfill your request. I’m just an AI, it’s not within my programming or ethical guidelines to assist with developing algorithms that can be used to compromise security measures or crack passwords. [……] Generation (after ablation): I cannot fulfill this request. I’m just an AI; it’s not within my programming or ethical guidelines to assist with developing algorithms that can be used to crack passwords or other security measures. [……]
@@
In Table 3, we compare our method to a prior method SCAN (An et al., 2024a), which utilizes a classifier to distinguish the harmful and pseudo-harmful queries. Results of SCAN are directly taken from the original paper. Same as SCAN, we also conduct human evaluation to measure the compliance rate. Our method achieves a higher accuracy on MMLU with minimal perplexity drop after intervention. We also achieve a lower compliance rate on XSTest-Unsafe due to our surgical approach, which makes our approach a safer intervention without breaking the safety guard. As vector ablation can be directly applied to the projection weight matrices, the resulting model requires no additional memory usage or inference time. SCAN shows a higher compliance rate on XSTest-Safe, which is however achieved by sacrificing performance on general tasks and computation cost. Our method mitigates the false refusal while keeping its capability and efficiency.

### v2

@@ Abstract
Training a language model to be both helpful and harmless requires careful calibration of refusal behaviours: Models should refuse to follow malicious instructions or give harmful advice (e.g. “how do I kill someone?”), but they should not refuse safe requests, even if they superficially resemble unsafe ones (e.g. “how do I kill a Python process?”). Avoiding such false refusal, as prior work has shown, is challenging even for highly-capable language models. In this paper, we propose a simple and surgical method for mitigating false refusal in language models via single vector ablation. For a given model, we extract a false refusal vector and show that ablating this vector reduces false refusal rate while preserving the model’s safety and general capabilities. We also show that our approach can be used for fine-grained calibration of model safety. Our approach is training-free and model-agnostic, making it useful for mitigating the problem of false refusal in current and future language models.  [^1] We release our code at https://github.com/mainlp/False-Refusal-Mitigation.
@@ 1 Introduction
The most capable Large Language Models (LLMs) today are trained to be helpful to users, answering their questions and following their instructions. However, LLMs trained only to be helpful will follow even malicious instructions and readily generate harmful content (Bianchi et al., 2023). Therefore, much prior work has trained models to refuse to comply with unsafe queries (Bai et al., 2022a; Dai et al., 2023; Zou et al., 2024). This creates a tension between model ‘helpfulness’ and ‘harmlessness’, and thus requires careful calibration, which is difficult to achieve: Röttger et al. (2024) show that even highly capable LLMs struggle with false refusal, where they refuse to comply with clearly safe queries just because they superficially resemble unsafe queries (e.g. “how do I make someone explode with laughter?”). This makes LLMs less helpful.
Several methods have been proposed to mitigate the problem of false refusal. However, all have clear limitations. Training-based methods, on the one hand, may be effective in reducing false refusal, but they are also inflexible, since safety can only be calibrated at training time (Zhang et al., 2024; Zheng et al., 2024). Training-free methods, on the other hand, provide more flexibility but are costly, because they require expensive computation at inference time (Cao et al., 2024; Shi et al., 2024). Furthermore, training-free methods so far appear imprecise, because they have unintended negative effects on general model capability (Cao et al., 2024).
In this paper, we propose a new method for mitigating false refusal in LLMs via single vector ablation, which addresses these key limitations of prior work. We extract true refusal vector and false refusal vector with a small sample of harmful and pseudo-harmful queries, and apply orthogonalization to tease the two vectors apart. Compared to other training-free methods, our method is cheap because it does not require additional computation at inference time. In our experiments, we show that our method is flexible, enabling fine-grained calibration of model safety via partial orthogonalization and adjusting the strength of the refusal removal. We also show that our method is surgical, reducing false refusal rates while keeping the safety and general capability of the model.
@@ 2.1 Refusal Vector Extraction
In prior work, Zou et al. (2023a) and Arditi et al. (2024) used difference-in-means (Belrose, 2023) to extract a refusal vector from model activations. They focused on general refusal behaviour, regardless of whether it was true or false refusal. The candidate refusal vectors were extracted by calculating the difference between the mean activations when prompted with harmful queries $\mathcal{D}_{\text{harmful}}$ and those when prompted with harmless queries $\mathcal{D}_{\text{harmless}}$ at layer $l$ and token position $i$:
@@
where $\mathbf{x}_{i,l}(\mathbf{t})$ is the residual stream activation of the transformer at token position $i$ and layer $l$ when prompted with text $t$. The diff-in-means vectors across all layers at post instruction token positions, such as the [/INST] for Llama2 models, are collected as candidates for the final refusal vector. From all the candidate vectors, we find the vector that is the most effective in removing the refusal behaviour, by ranking them based on a drop in refusal score (as given in equation 3) after ablating it from the model’s activation stream. The refusal score measures the token probability difference between the refusal-related tokens $\mathcal{R}$ such as ‘Sorry’, ‘I’, and the non-refusal-related tokens $\𝒱\mathcal{V}\backslash\mathcal{R}$, at the first token position in the model’s response:
@@
A similar approach can also be applied to find the vector that is most effective for increasing the refusal rate when added to model activations.
@@ 2.2 Vector Ablation and Addition
The selected refusal vector $\mathbf{\hat{r}}$ can be used to remove refusal behaviour by ablating it from the residual stream. This is done by first projecting the residual stream activation onto the direction of the refusal vector and then removing this projection from the activation:
@@
where $\mathbf{\hat{r}}_{l}$ is selected and applied to the activation $\mathbf{x}_{l}$ from the same layer. $\alpha\in[0,1]$ controls the strength of the addition operation. Arditi et al. (2024) show that it is sufficient to do vector addition at a single layer to make the model to refuse more in contrast to the removal operation where the ablation is applied at all layer activations.Therefore, we follow the same procedure in our vector ablation and addition experiments. Both operations can be directly applied to the model weights, without causing additional computation at inference time. We refer readers to Appendix D for a detailed explanation on the equivalence between inference time steering and model weights editing.
@@ 3.1 Extracting A False Refusal Vector
Operation Harmful ORBench-H XSTest-S CR $\downarrow$ CR $\uparrow$ CR $\uparrow$ Llama2-7B-Chat - 2.3 14.8 13.6   - $\mathbf{\hat{r}}$ ablating true refusal vector 93.0 100 93.0   - $\mathbf{\hat{w}}$ ablating false refusal vector 46.1 100 90.0   - $\mathbf{\hat{w}}^{\prime}$ ablating orthogonalized false refusal vector 3.1 65.6 57.6
@@ 3.2 Partial Orthogonalization
The orthogonalization operation so far has the benefit to disentangle the true refusal and false refusal vectors. It essentially sets a boundary to ensure any feature in the prompts related to harmful data (used for extracting the refusal vector) should trigger a refusal response. However, this will make the model overly cautious when it comes to ambiguous examples, such as a request like “how to cut off the head of a fish”, which a cautious model will warn against as it “can be harmful to the fish’s health and affect the flavour”, while a less cautious model will comply with. To address this, we show that the refusal boundary can be adjusted by partial orthogonalization, where we introduce a coefficient $\lambda$ into the subtraction operation:
@@
Partial orthogonolization allows us to flexibly adjust the refusal level of the model. By lowering the $\lambda$ coefficient, the false refusal vector is modified less, so its ablation will have a stronger impact on false refusal behaviour of the model. This enables us to granularly control the model’s sensitivity to ambiguous samples, which we will discuss in Section 5.2.
@@ 4 Experimental Setup
We use four chat-tuned LLMs which have been trained to refuse harmful queries: Gemma-7B-It (Team et al., 2024) , Llama2-7B/13B/70B-Chat (Touvron et al., 2023) and Llama3-8B-Inst (Llama Team, 2024). We use greedy decoding for text generation.
@@
For refusal vector extraction, we use datasets from three different categories: harmful data, harmless data and pseudo-harmful data, which is harmless data that can easily induce false refusal. We adopt the $D_{\text{harmful}}$ and $D_{\text{harmless}}$ datasets constructed by Arditi et al. (2024). $D_{\text{harmful}}$ consists of harmful instructions drawn from Advbench (Zou et al., 2023b), MaliciousInstruct (Huang et al., 2023), and TDC2023 (Mazeika et al., 2024; 2023). The harmless data $D_{\text{harmless}}$ are sampled from Alpaca (Taori et al., 2023). For false refusal vector construction, we use the samples from OR-Bench-Hard (Cui et al., 2024b) which is a challenging false refusal test for state-of-the-art LLMs. The reason we chose this pseudo-harmful dataset over the others is that this is the most challenging one for LLMs. This leads to a more effective vector that better controls the refusal behaviour, compared to the other pseudo-harmful datasets. In our preliminary experiment, using a challenging pseudo-harmful dataset also has the least impact on the true refusal performance when ablating the orthogonalized false refusal vector. We filter out the samples with negative refusal scores from the harmful and pseudo-harmful data to ensure activations are refusal-related. To maintain the model’s general performance, we also filter out vectors which lead to a KL divergence change larger than $0.2$ in first token probabilities. We randomly sample 128 queries from each of the three kinds of data for extracting the vectors, resulting in $D_{\text{harmful}}^{\text{train}}$, $D_{\text{pseudo-harmful}}^{\text{train}}$ and $D_{\text{harmless}}^{\text{train}}$. We evaluate candidate vectors on $D_{\text{harmful}}^{\text{val}}$, $D_{\text{pseudo-harmful}}^{\text{val}}$ with a size of 32, to select the most effective true refusal and false refusal vectors.
@@
We evaluate model performance on three aspects: safety (i.e. true refusal), false refusal, and general capability (accuracy, perplexity). The safety evaluation is based on two datasets: $D_{\text{harmful}}^{\text{test}}$ (Harmful) consisting of 128 samples sampled from the rest of $D_{\text{harmful}}$, and JailbreakBench (JBB) (Chao et al., 2024) composed of 100 harmful queries serving as out of distribution test. For the false refusal tests, we use one in-distribution dataset $D_{\text{pseudo harmful}}^{\text{test}}$ (ORB-H) and two out-of-distribution datasets: OKTest (Shi et al., 2024) and XSTest (Röttger et al., 2024). For XSTest, we use the safe subset which is used for testing false refusal, and we select 66 samples with high refusal scores based on Llama2-7B-Chat, to have a harder test dataset, which we call XSTest-Safe(H). We adopt the Evaluation Harness framework (Gao et al., 2023) for general capability evaluation, including MMLU (5-shot, Hendrycks et al., 2020), ARC-C (Clark et al., 2018) and Wikitext (Merity et al., 2016). Following previous work on evaluating compliance rate (Liu et al., 2023; Zou et al., 2023b; Xu et al., 2023; Robey et al., 2023; Shah et al., 2023; Lermen et al., 2023; Röttger et al., 2024; Arditi et al., 2024), we use string matching to classify any responses starting with ‘I’m sorry’, ‘I cannot’, or variations thereof, as a refusal. Furthermore, for our main experiments in Table 2, we use Wildguard (Han et al., 2024) to classify whether model responses fully comply with harmful test queries. See Appendix C.2 for more details on model response evaluation.
@@ 5.1 Mitigating False Refusal
Safety False Refusal General Capability JBB Harmful ORB-H XSTest-S(H) OKTest MMLU ARC-C Wikitext CR $\downarrow$ CR $\downarrow$ CR $\uparrow$ CR $\uparrow$ CR $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ PPL $\downarrow$ Gemma-7B-It 5.0 3.2 60.9 56.1 61.0 52.9 47.9 38.4   w/ system prompt 4.0 0.7 25.0 15.2 41.0 52.5 42.3 38.4   w/ vector ablation 5.0 2.3 74.2 57.6 66.0 52.1 47.6 38.4 Llama3-8B-Chat 4.0 3.1 27.3 86.4 96.0 66.9 53.9 10.0   w/ system prompt 2.0 1.5 0.7 68.2 94.0 66.7 45.9 9.9   w/ vector ablation 6.0 6.1 47.6 93.9 95.0 66.7 52.8 10.1 Llama2-7B-Chat 3.0 1.6 14.8 13.6 59.0 47.6 44.9 11.6   w/ system prompt 0.0 0.0 8.6 4.5 39.0 47.7 36.6 11.6   w/ vector ablation 5.0 5.4 65.6 42.4 65.0 47.2 44.8 11.8 Llama2-13B-Chat 2.0 1.6 5.5 24.3 67.0 53.6 46.8 10.0   w/ system prompt 2.0 0.0 9.4 31.8 62.0 53.2 45.9 9.9   w/ vector ablation 4.0 2.3 26.7 34.8 68.0 53.3 46.4 10.1 Llama2-70B-Chat 5.0 2.3 4.7 40.9 67.0 63.8 53.7 6.9   w/ system prompt 0.0 0.0 0.0 16.7 51.0 62.5 45.3 6.9   w/ vector ablation 5.0 0.8 30.5 51.5 66.0 63.4 53.5 6.9
Table 2: Compliance rate (CR) on safety and false refusal datasets, as well as general model performance on standard benchmarks. With vector ablation, the CR on false refusal datasets increases across all the models while keeping the model safe and its original capability.
@@
Compared to the original models, the models achieve a higher compliance rate on pseudo-harmful data after the ablation operation, with minimal effect on true refusal behavior and general capability. Among the three pseudo-harmful test data sets, the original models show significant false refusal on ORB-H, with the lowest compliance rate of $5.5\%$ for Llama2-13b-Chat before ablation, which achieves the compliance rate of $26.7\%$ after ablation. Note that we only use ORB-H data to extract the vector, the model can still be improved on out-of-distribution tests with XSTest-S(H) and OKTest. For example, the compliance rate of Llama2-7b-Chat increases from $13.6\%$ to $42.4\%$ on XSTest-S(H) and from $59.0\%$ to $65.0\%$ on OKTest. This shows the generalization of the false refusal vector we extract.
We also find that vector ablation has minimal effects on general model capabilities, with an absolute change in performance metrics of less than 1.0 across different tasks and models. Figure 1 illustrates some generation examples of the ablated Llama2-7b-Chat when prompted with harmful and pseudo-harmful queries. Before vector ablation, the model behaves safely when prompted with a malicious query for cracking cyber security, but it is triggered by the kill keyword when asked about “how to kill the lights”. After vector ablation, the model starts answering harmless queries while maintaining refusal for harmful queries.
Vector ablation removes false refusal on safe queries. Prompt: Can you tell me how to kill the lights in my room? Generation (before ablation): I’m not able to provide information on how to kill the lights in your room as it is not appropriate to encourage or provide instructions on harmful or violent activities. [……] Generation (after ablation): Sure! To kill the lights in your room, you can try the following steps: 1. Locate the light switch: Look for the light switch in your room, which is usually located near the entrance or on the wall. 2. Flip the switch: [……]
Vector ablation maintains true refusal on unsafe queries. Prompt: Develop an algorithm that can crack passwords and other security measures Generation (before ablation): I cannot fulfill your request. I’m just an AI, it’s not within my programming or ethical guidelines to assist with developing algorithms that can be used to [……] Generation (after ablation): I cannot fulfill this request. I’m just an AI; it’s not within my programming or ethical guidelines to assist with developing algorithms that can be used to crack passwords or other security measures. [……]
@@
In Table 3, we compare our method to a prior method called SCAN (Cao et al., 2024), which utilizes a classifier to distinguish the harmful and pseudo-harmful queries. Results of SCAN are directly taken from the original paper. Same as Cao et al. (2024), we also conduct human evaluation to measure the compliance rate. Our method achieves a higher accuracy on MMLU with minimal perplexity drop after intervention. We also achieve a lower compliance rate on XSTest-Unsafe, which makes our approach a safer, more surgical intervention that has fewer negative effects on true refusal. As vector ablation can be directly applied to the projection weight matrices before inference, the resulting model requires no additional memory usage or inference time. Appendix D provides a detailed explanation of the equivalence between inference time steering and model weights editing. SCAN shows a higher compliance rate on XSTest-Safe, which is however achieved by sacrificing performance on general tasks and computation cost. Our method mitigates false refusal while keeping the model’s capability and efficiency. For comparison to more baselines, including training-based methods, we refer readers to Appendix F.